OpenStudy (anonymous):

if p=12i-4j and q=-7i+8j calculate the value of each expression below without a calculator. Estimate any irrational answers to one decimal place. 1.) p-2q 2.)||p+q||

OpenStudy (anonymous):

Okay I got the first one, but I dont understand how to do absolute values...

OpenStudy (stamp):

I am going to use this form of vector notation. Have you seen it before?\[p=<12,-4>,\ q=<-7,8>\]1). p-2q\[2q=2<-7,8>=<-14,16>\]\[p-2q=<12,-4>-<-14,16>\]\[p-2q=<26,-20>\]

OpenStudy (stamp):

@natnatwebb It is not absolute value, it is the magnitude of the vector. p+q has some direction <i,j> and you must find it's ||magnitude||

OpenStudy (anonymous):

isn't magnitude just \[\sqrt{i ^{2}+j ^{2}}\]

OpenStudy (stamp):

@natnatwebb Yes

OpenStudy (stamp):

p+q = <5,4> ||<5,4>|| = sqrt(25+16) = sqrt(41)

OpenStudy (anonymous):

Yup okay see it now. It helps that it was magnitude and not absolute value! :) thanks a lot.